OpenStudy (anonymous):

Solve the given linear-quadratic system graphically. y = −(x − 2)2 + 3 y = −4 Please lists steps... I need to know how to do it

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

You would need to graph both equations to solve. Are you able to do so?

OpenStudy (anonymous):

set \[-(x+2)^2+3=-4\] and solve for \(x\)

OpenStudy (anonymous):

@satellite73 graphically no algebraically

OpenStudy (anonymous):

oops typo there sorry \[-(x-2)^2+3=-4\] solve for \(x\) takes three steps

OpenStudy (anonymous):

The only thing that is bothering me is the - before the "(" How would I foil it?

OpenStudy (anonymous):

oh no!!!

OpenStudy (anonymous):

don't foil nothing

OpenStudy (anonymous):

lets go through the steps, there aer only three of them

OpenStudy (anonymous):

Lol. How would I do it then. I don't know how to do it graphically or algebraically

OpenStudy (anonymous):

Are you able to graph y=-4?

OpenStudy (anonymous):

\[-(x-2)^2+3=-4\] subtract 3 from both sides what do you get?

OpenStudy (anonymous):

wouldnt you move four over?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

-(x-2)^2+7

OpenStudy (anonymous):

You need to isolate the variable.

OpenStudy (anonymous):

subtract 3 from both sides of the equal sign

OpenStudy (anonymous):

don't add 4 to both sides,you would be working backwards

OpenStudy (anonymous):

@satellite73 He won't understand if you are just giving directions, you must explain why.

OpenStudy (anonymous):

so... -(x-2)^2=-7

OpenStudy (anonymous):

right now change the sing of both sides, i.e. get rid of both minus signs

OpenStudy (anonymous):

*sign

OpenStudy (anonymous):

(x-2)^2=7

OpenStudy (anonymous):

explanation comes with the doing

OpenStudy (anonymous):

right

OpenStudy (anonymous):

now take the square root of both sides, i.e. get rid of the square on the left don't forget the \(\pm\)

OpenStudy (anonymous):

x=2\[\pm \sqrt{7}\]

OpenStudy (anonymous):

that sort of came out right

OpenStudy (anonymous):

bingo

OpenStudy (anonymous):

\[x=2\pm\sqrt7\]

OpenStudy (anonymous):

here is the graph if you really need it http://www.wolframalpha.com/input/?i=y%3D-%28x-2%29^2%2B3%2Cy%3D-4

OpenStudy (anonymous):

x=4.6 or -.645

OpenStudy (anonymous):

so then I plug it in to get Y.

OpenStudy (anonymous):

notice no foil was harmed in solving this question

OpenStudy (anonymous):

lol no you don't plug it in to get \(y\) you are told at the outset that \(y=-4\)

OpenStudy (anonymous):

Oh yea, thats true

OpenStudy (anonymous):

The solution to the system is (−1, −6) and (5, −6). thats the answer

OpenStudy (anonymous):

@satellite73 would I round the answer to 5?

OpenStudy (anonymous):

@satellite73 Anyway you can help me?

OpenStudy (anonymous):

i am lost now

OpenStudy (anonymous):

unless perhaps there is a typo in the question

OpenStudy (anonymous):

@Beleaguer It would've been easier if you graphed the two.

OpenStudy (anonymous):

could it have actually been \[-(x-2)^2-3=-4\]?

OpenStudy (anonymous):

thats exactly how I am right now

OpenStudy (anonymous):

Just graph the two and see where they meet.

OpenStudy (anonymous):

actually it makes no sense anyway how can the answer be (−1, −6) and (5, −6). when you are told \(y=-4\) was it really \[y=-6\]?

OpenStudy (anonymous):

Nope, its +3 = -4. I pasted it from the book. Maybe the book is wrong? lol i have no idea

OpenStudy (anonymous):

now that i see it, it is either a typo on their part of yours

OpenStudy (anonymous):

\[-(x-2)^2+3=-9\\ -(x-2)^2=-9\\ (x-2)^2=-9\\ x-2=3, x-2=-3\\ x=-1,x=-5\]

OpenStudy (anonymous):

now i made a typo, first line should have been \[-(x-2)^2+3=-6\]

OpenStudy (anonymous):

if they wrote \(y=-4\) they meant \(y=-6\) that is clear, because both answers have \(-6\) in the second coordinate

OpenStudy (anonymous):

What about this y = −(x − 2)2 + 8 y = −9

OpenStudy (anonymous):

I got x= 6.123

OpenStudy (anonymous):

lorda mercy lets assume there is not typo here stop with the decimals already

OpenStudy (anonymous):

thats what we are supposed to use

OpenStudy (anonymous):

x= 2+rad17

OpenStudy (anonymous):

\[--(x-2)^2+8=-9\\ -(x-2)^2=-17\\ (x-2)^2=17\\ x-2=\pm\sqrt{17}\\ x=2\pm\sqrt{17}\]

OpenStudy (anonymous):

yeah you got it

OpenStudy (anonymous):

thats what I got, but woah...theres more xD

OpenStudy (anonymous):

this is the answer -.- The solution to the system is (−1, −1) and (5, −1).

OpenStudy (anonymous):

I am going to look on the internet on how to solve it graphically..

OpenStudy (anonymous):

the formula for vertex is -b/2a, right?

OpenStudy (anonymous):

can you post a screen shot?

OpenStudy (anonymous):

if \(y=-9\) it cannot be \(-1\) at the same time

OpenStudy (anonymous):

that is the answer if you solve \[-(x-2)^2+8=-1\]

OpenStudy (anonymous):

either there are massive typos here, of there is a problem in the translation

OpenStudy (anonymous):

you cannot have \(y=-9\) and have a solution that has anything other than \(-9\) in the second coordinate maybe your math teacher was on the sauce when he/she wrote these, or copied and pasted incorrectly

OpenStudy (anonymous):

\[-(x-2)^2+8=-1\\ -(x-2)^2-9\\ (x-2)^2=9\\ x-2=3,x-2=-3\\ x=5,x=-1\] solutions are \((5,-1),(-1,-1)\)

OpenStudy (anonymous):

but that is \(y=-1\) NOT \(y=-9\)

OpenStudy (anonymous):

I guess it is because we are solving it graphically can you help me that way @satellite73

OpenStudy (anonymous):

it is the same graphically!!

OpenStudy (anonymous):

solutions will be the same, makes no difference

OpenStudy (anonymous):

Well, I literally pasted exactly what I saw

OpenStudy (anonymous):

there is the graphical solution to \[y=-(x-2)^2+8\\ y=-1\] notice that it is \((5,-1),(-1,-1)\) http://www.wolframalpha.com/input/?i=y%3D-%28x-2%29^2%2B8%2Cy%3D-1

OpenStudy (anonymous):

i know you are confused because of the mistake in the question, but try to ignore that if \(y=-1\) then \(y=-1\) and if \(y=-9\) then \(y=-9\) it cannot be otherwise

OpenStudy (anonymous):

Ok, lets try something else.... maybe I will understand it. How about this problem. \[y=-(x-1)^{2}+11 y= x+4\]

OpenStudy (anonymous):

perhaps the teacher made a mistake and did the first step mentally and therefore wrote \[-(x-2)^2+8=-9\] instead of \(-(x-2)^2+8=-1\) because the first step in solving \[-(x-2)^2+8=-1\] would be to turn it in to \[-(x-2)^2=-9\]

OpenStudy (anonymous):

First step is set it equal to each other?

OpenStudy (anonymous):

it is not a matter of you understanding there is a mistake in the question, take my word for it

OpenStudy (anonymous):

Well this is from another source, so I am going to attempt it and see

OpenStudy (anonymous):

I don't recall learning the other way

OpenStudy (anonymous):

look if you went to solve a simple one like \[y=2x+1,y=9\] you would set \[2x+1=9\\2x=8\\ x=4\] and the solution would be \((4,9)\) not some other number in the second slot

OpenStudy (anonymous):

but what about the one I just posted

OpenStudy (anonymous):

\[-(x+1)^2=11\\ y=x+4\] is a different kind of question because of the \(x+4\) in this case you would have to solve \[-(x+1)^2+11=x+4\] here you don't know what \(y\) is, so you would have to find it

OpenStudy (anonymous):

Where do I go from there

OpenStudy (anonymous):

for this one, you need to put it all on one side of the equal sign and solve \[(x+1)^2+x+4-11=0\]

OpenStudy (anonymous):

So by moving the 4 over the - goes away?

OpenStudy (anonymous):

then expand and combine like terms \[x^2+2x+1+x-7\\ x^2+3x+6=0\]

OpenStudy (anonymous):

i added \((x+1)^2\) and subtracted \(11\) from both sides easier to work with \((x+1)^2\) than \(-(x+1)^2\)

OpenStudy (anonymous):

any chance i can convince you to post a screen shot of the original two questions?

OpenStudy (anonymous):

Yea, I will. I just want to get this problem straight because I know this will be on my test tomorrow 100%, im not sure about the other ones

OpenStudy (anonymous):

hope your teacher doesn't make the same mistake on the test!

OpenStudy (anonymous):

So, do you mind going over the process of eliminating the -? I am lost on that part, sorry for being so picky and ty for sticking through

OpenStudy (anonymous):

let me make sure i have the question correct is it \[y=-(x+1)^2+11\\ y=x+4\]?

OpenStudy (anonymous):

I keep thinking I would distribute the - into the x and -1 therefore making it (-x+1)^2

OpenStudy (anonymous):

hell no!!

OpenStudy (anonymous):

Yes